OpenStudy (anonymous):

Select all that apply. Select the formulas that give the amount P in the account after the first period. A. P= 1.005Po B. P= 1.02Po C. P= Po(0.02/4) D. P=Po(1.02/4) F. P=Po(1+ 0.02/4)

OpenStudy (anonymous):

is there more to this question or no?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

https://hanwingspanps.owschools.com/media/g_alg02_ccss_2015/9/img_g_alg02u09p26d_2a.gif this is what one of the equations actually looks like though

OpenStudy (welshfella):

theres not enough information given to answer this question

OpenStudy (anonymous):

Time period Amount in the account (with rounding) Initial principal, P 0 1,000 End of the first period 1,000+1,000=0.02/4=1005

OpenStudy (welshfella):

that looks like compound interest at 2% paid 4 times in the first year

OpenStudy (anonymous):

1,000+1,000+0.02/4=1005

OpenStudy (anonymous):

that looks correct

OpenStudy (welshfella):

P = Po(1 + 0.02/4)

OpenStudy (welshfella):

thats option F

OpenStudy (welshfella):

theres another one which applies also Can you spot it?

OpenStudy (anonymous):

there is no option f

OpenStudy (welshfella):

the one after D then

OpenStudy (welshfella):

the last one but theres another whwich will give you the same result 1005.

OpenStudy (anonymous):

its supposed to be e sorry

OpenStudy (anonymous):

probably d

OpenStudy (welshfella):

No - my mistake - i thought d but thats not correct.

OpenStudy (welshfella):

its just the last one.